OpenStudy (anonymous):

Eve can swim 3 m/s in still water. The river current is 1 m/s. The river is 200 m wide. Eve wishes to swim across the river, 20 m upstream, to meet Adam underneath the apple tree. a) In what direction should Eve swim (relative to the riverbank)? b) What will be her resultant speed? c) How long will Adam have to wait under the apple tree?

OpenStudy (anonymous):

ooh physics. i think you have to use vector formulas such as \[V _{x}=v \cos \theta\] and \[V _{y}=v \sin \theta\]

OpenStudy (anonymous):

and trig functions

OpenStudy (anonymous):

I still don't get it. /:

OpenStudy (anonymous):

Oh yes. This again. I will leave you to the more than capable people here =P